Sharps Fire


CONTAINED

Incident #: EIS-247
Fire Start: 3 months ago
Location: 4.5 miles NE of Bellevue, ID
Acres: 64,763.5
Agency: South Central Idaho Interagency Dispatch Center (ID-SCC)
Dispatch Notes: Contain: 9/6 @ 1026control: 9/11 @ 1700

Current Weather

Fire Growth Potential

Incident Overview

Updated: 4 mins, 41 secs ago — FINAL UPDATE 8.16.18 The Sharps Fire is transiting from a type 3 incident management team to a local type 4 incident. Resources will remain on scene monitoring the fire and mopping up hotspots. At this time, the road and trail closures will remain in effect. Road and trail closures remain in effect. Closed roads on the BLM include Upper Quigley Canyon and Cove Creek Roads. Closed roads on the Sawtooth National Forest include Baugh Creek #115A and Copper Creek #134. The following trails on the Sawtooth National Forest are also closed: Fisher Canyon, Porcupine Creek Kale Creek, Federal Gulch, Timber Draw and Iron Mine Creek (refer to Map).For updated road and trail closures please visit:Bureau of Land Management Road Closures: click here Sawtooth National Forest Road and Trail Closures: click here Weather continues to be hot and dry in southern and central Idaho. Fire Restrictions Remain in EffectFor additional updates on the Sharps Fire visit Idaho Fire Info Facebook or Sawtooth...

Contact

Incident Contact
s Sharps Fire
Email: 2018.sharpsfire@firenet.gov
Phone: 208-731-8604
Hours: 7:00 am - 9:00 pm Daily Julie Thomas
Phone: 208-731-8604Show additional contacts
Fire Unit Information
Twin Falls District
Bureau of Land Management
400 West F Street
Shoshone, ID 83352

Basic Information

Cause
Human-caused
Date of Origin
Sunday July 29th, 2018 approx. 11:15 AM
Location
6 miles east of Bellevue Idaho
Incident Commander
Type 4 Incident
Coordinates
43.467 latitude, -114.145 longitude

Current Situation

Size
64,853 Acres
% Contained
90%
Estimated Containment Date
Friday September 21st, 2018 approx. 12:00 AM
Fuels Involved
Timber, short grass, brush
Significant Events
Today the fire had minimal fire behavior with creeping and smoldering.

Outlook

Planned Actions
Hand crews are continuing to cold trail and secure containment lines. Personnel are also mopping up interior hear pockets that are appearing in timber stringers and duff pockets due to the high temperatures. Crews on the west and northwest portions of the fire continue implementing suppression repair utilizing heavy equipment and hand crews. Suppression repair on the southern end of the fire is complete. READ's are assessing repair needs and working with suppression repair personnel on Forest Service lands.
Projected Incident Activity
Low fire behavior and activity expected.